"Compute the depreciation each year of a machine that costs $68,000 to purchase and $3,100 to install with 10 years of life. Use DDB method (multiplier = 2) switching to straight line depreciation when appropriate. Assume a salvage value of $0. How much should be depreciated in year 8?"
Answer #1
Purchase Cost $68,000
Installation Cost $3,100
Total Cost $71,100
Year DDB Method Remaining value Straight line method
1 71,100*20% = 14,220 71,100 - 14,220 = 56,880 71,100/10 = 7,110
2 56,880*20% = 11,376 56,880 -11376 = 45504 56,880/9 = 6,320
3 45,504*20% = 9,100.8 45,504 - 9100.8 = 36403.2 45,504/8 = 5,688
4 36403.2 *20% = 7,280.6 36,403.2- 7,280.6 = 29122.6 36,403.2/7 = 5,200.5
5 29,122.6*20% = 5,824.5 29,122.6 - 5,824.5 = 23,298 23,298/6 = 3,883
6 23,298*20% = 4,659.6 23,298 - 4,659.6 = 18,638 23,298/5 = 4,659.6
In year 6 the depreciation value is same for both method and in 7th year DDB depreciation value is lesser than SLM depreciation method hence we switched to SLM method from year7. At year 8 depreciation value is $ 4,659.5
Depreciation using SLM in year 7 18,638/4 = 4,659.5
Remaining value
7 18,638*20% = 3,727.7 18,638 - 3,727.7 = 14910.8 18,638.4/4 = 4,659.6 18,638 - 4,659.5 = 13978.4
8 14910.8 *20% =2,982.1 14910.8 - 2,982.1 = 11928.6 14,910.8/3 = 4,970.3 13,978 - 4,659.5 = 9,318.9
9 11,928.6*20% = 2,385.7 11,928.6 - 2,385.7 = 9542.9 11,928.6/2 = 5,964.3 9,318.9 - 4,659.5 = 4,659.4
10 9,542.9*20% = 1,908.6 9,542.9 - 1,1908 =7634.3 9,452.9/1 = 9,452.9 4,659.4- 4,659.5 = 0
Depreciation applied for each year